Project

General

Profile

Feature #1854

Muon Beamline Step IV prep list

Added by Nebrensky, Henry almost 7 years ago. Updated over 6 years ago.

Status:
Closed
Priority:
Normal
Category:
Beam Line
Start date:
06 June 2016
Due date:
28 June 2016
% Done:

100%

Estimated time:

Description

List of things to sort out by start of Step IV - end June 2016


Related issues

Follows Operations - Bug #1788: Muon Beamline Step IV Prep listClosedNebrensky, Henry28 October 201516 February 2016

Actions
Precedes Operations - Bug #1876: Muon Beamline Step IV prep listClosedNebrensky, Henry21 September 201607 November 2016

Actions
#1

Updated by Nebrensky, Henry almost 7 years ago

Stuff that needs sorting out by 28th June 2016.
Important stuff in bold

Target

  • Light Loss (TGT) - stabilised in mild saturation again
  • Trace fibre connections - January shutdown awaiting MICE Note
  • Restart loopback measurements (TGT)
  • Protocol for synch access for swapping fibres at patch panel (TGT)
  • Return/replace target1ctl
  • S3 and T2 - assembly of next spare (TGT) - pinch point is getting optics block off T1
  • Clean-up of R78 target control crate (DL) need to restart VME control (and possible move to R9_
  • ISIS Target - cooling water flow read-out?
  • Target Alarm handler #1613 update 22 (CAM)
  • ISIS data items #1669 (CAM)
  • Telephone link #977
  • Activation Study (activation seen: limit of 2 V at MS/64 stands)
  • Beambump tune
  • Updated target manual (EO/JL)

Conventional Magnets

  • review existing water flows and temperature range (MBC)
    • need max current limits for quads in control software? (CAM)
  • Calibrate D1 & D2 current reading and update IOC software (DL)
  • Q4 oscillation below 10A and anomalous ramp rate (DL) reg board change
  • 50 A minimum current limits for D1, D2 in control software (CAM)
  • Q1 MPS start-up recalcitrance (DL) MPS seems happy in new location
  • Diagnose flooded Danfysik 9000 supply
    • faulty PCB - Richard Hale contacting Danfysik about a replacement
    • Have ordered for replacement board
  • Q1 alarm on -ve voltage at zero amps (CAM)
    • Now on both Q1 and Q2 MPS
  • Q4 Earth Leakage (MBC) - D2,Q4,Q5 flushed through
    • Q4 threshold loosened
  • Q5 leaks fitting tightened, the other has disappeared
  • Q6 leak leave until time
  • Leak sensors and drip trays
  • "All Off" controls misbehaviour
  • Water, temperatures and MPS water flows/temperatures in ALH/archiver (CAM)
    • *Q6 temperature alarm limits missing!*
    • Interlocks alarm is inoperative
  • Update magnet and POMPOM user guides (MBC)
  • Annual POMPOM radiation damage test (MBC)
    • negative
    • positive

Decay Solenoid

  • Overheating investigation (DK/DL)
  • Replace PSU (DK/DL)
    • Calibrate current reading and update IOC software (DL)
  • 120A trip investigation (DK/DL)
  • Replace final battery cell (DL - confirmed still outstanding )
  • Diodes in BBU circuit (DL)
  • Boosted air (MBC)
    • Trench sensor (awaiting read-out)
    • Compressor sensor
    • Backup compressor
  • Linde PC install: awaiting validation and IOC (DK) - licence issues
    • MLCR monitor and KVM implications (HN)
  • ALH/archiver corrections #1710 (CAM)
  • Magnet-On status lamp (DL) - needs synch access

Proton Absorber

  • Combine survey documents (MBC)
  • 15mm slab - not reproduced

Beamstop

  • Translate Survey document (MBC)

General

  • BLOC training and experience (MBC)
  • Shared services/environment (water, air, computing/networking, UPS ( #1213 ) in Alarm Handler and in Archiver (CAM)

Beyond the Muon Beamline

#2

Updated by Nebrensky, Henry over 6 years ago

  • Status changed from Open to Closed
  • % Done changed from 0 to 100

Pass baton to #1876

Also available in: Atom PDF